OpenStudy (anonymous):

solve 1 + cosx over 1- cosx = 0

OpenStudy (ash2326):

we have \[\frac{1+\cos x}{1-\cos x}=0\] 1- cos x can't be infinity so \[1+cos x=0\] or \[\cos x=-1\] so x=180, 540 , 900 and so on

OpenStudy (anonymous):

What do you me by infinity?

OpenStudy (ash2326):

Buster kitten infinity is a very large number, you know tan 90 is infinity. It's defined as \[\frac{any\ number}{0}\] so \[\frac{1}{infinity}=0\]

OpenStudy (ash2326):

@Busterkitten I'm sorry. Kept you waiting

OpenStudy (anonymous):

nah, it's all good and thank you. Sorry if I may not be the sharpest tool in the shed.

OpenStudy (ash2326):

@Busterkitten I feel you're a smart girl:)

OpenStudy (anonymous):

Thanks, you seem very knowledgeable with math :D. Thanks for all of your help :D